# THE RESTRICTED PICARD FUNCTOR

### v2

@@ 6 Structure of the restricted Picard functor
We may assume that $K=K_{s}$ and that $X$ is nonempty and connected. Because $X$ is smooth, $X(K)\neq\emptyset$, so Proposition 6.6 implies that $\operatorname{Pic}(X\times S)\twoheadrightarrow\operatorname{Pic}^{+}_{X/K}(S)$ is surjective for every smooth $S/K$. We need to show that there are no nonconstant maps to $\operatorname{Pic}^{+}_{X/K}$ for $S$ unirational. We may assume that $S$ is open in $\mathbf{P}^{1}$. Then $\operatorname{Pic}^{+}_{X/K}(S)=\operatorname{Pic}(X\times S)=\operatorname{Pic}(X)$, the last equality by [EGAIV4, Cor. 21.4.11]. Thus, any map $S\rightarrow\operatorname{Pic}^{+}_{X/K}$ is constant – that is, is obtained via scalar extension from a map ${\rm{Spec}}(K)\rightarrow\operatorname{Pic}^{+}_{X/K}$. ∎
Our next goal is to obtain representability of $\operatorname{Pic}^{+}_{X/K}$ by a scheme when every map from $X$ into an abelian variety has finite image (Theorem 1.4).
Theorem 6.8.
((((Theorem 1.4)))) Let $K$ be a field and $X$ a regular $K$-scheme of finite type with dense smooth locus. Assume that any map from $X_{K_{s}}$ into an abelian variety $A/K_{s}$ factors through a finite subscheme of $A$. Then $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth commutative $K$-group scheme with finitely generated component group. Furthermore, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is wound unipotent.
Proof.
Let $\operatorname{Pic}^{+}_{X/K}\simeq G/E$ be an almost-representable presentation of $\operatorname{Pic}^{+}_{X/K}$ (Theorem 5.3). By Corollary 6.5, $G^{0}$ is an extension of a smooth unipotent group scheme by an abelian variety $A$. The key point is to show that $A=0$. We may assume that $K$ is separably closed and that $X$ is nonempty and connected, hence irreducible. The map $A\rightarrow G/E\simeq\operatorname{Pic}^{+}_{X/K}$ yields an element $\alpha\in\operatorname{Pic}^{+}_{X/K}(A)$. By Corollary 6.2, the obstacle to $\alpha$ arising from an element of $\operatorname{Pic}(X\times A)$ lives in ${\rm{H}}^{2}(A,\mathbf{Z}^{n})$ for some $n\geq 0$. By Lemma 6.3, we have a functorial isomorphism ${\rm{H}}^{2}(A,\mathbf{Z})\simeq{\rm{H}}^{1}(A,\mathbf{Q}/\mathbf{Z})$. The latter agrees with ${\rm{Hom}}_{\rm{cts}}(\pi_{1}(A),\mathbf{Q}/\mathbf{Z})$. Any connected finite étale cover of $A$ is itself an abelian variety. We thus obtain an abelian variety $B$ – nonzero if $A$ is – and a finite homomorphism $B\rightarrow G^{0}$ such that the induced element of $\operatorname{Pic}^{+}_{X/K}(B)$ comes from a line bundle $\mathscr{L}$ on $X\times B$.
Now $\mathscr{L}$ defines a map $X\rightarrow\widehat{B}$ to the dual abelian variety. By hypothesis, this map has finite image. Because $X$ is integral and $X(K)\neq\emptyset$, it is a constant map to a point $b\in B(K)$. This means that $\mathscr{L}\in\operatorname{Pic}(X)\oplus\operatorname{Pic}(B)$, so the map $B\rightarrow\operatorname{Pic}^{+}_{X/K}$ is constant. This in turn means that the map $h\colon B\rightarrow G$ factors through $E\subset G$. Because $B$ is connected and $E$tale, it follows that $h=0$, hence $A=0$. By Corollary 6.5, it follows that $G^{0}$ is smooth and unipotent, and even trivial if $K$ is perfect. Thus $G^{t}$ is torsion. It follows that $E\subset G^{t}$ is trivial. Thus $\operatorname{Pic}^{+}_{X/K}\simeq G$ is a smooth $K$-scheme. Finally, $(\operatorname{Pic}^{+}_{X/K})^{0}$ is wound unipotent by Proposition 6.4. ∎
@@
The second statement of the theorem follows from Proposition 6.6, because $\mathbf{A}^{n}$ has no non-constant units over a field. Because $X$ is smooth, $X(K_{s})$ admits a $K_{s}$-point. Because $\mathbf{A}^{n}$ has no non-constant map to an abelian variety, Theorem 6.8 implies that $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth $K$-group scheme with finitely-generated component group. Proposition 6.7 implies that $(\operatorname{Pic}^{+}_{X/K})^{0}$ is strongly wound, so it only remains to show that the component group of $\operatorname{Pic}^{+}_{X/K}$ is $p$-power torsion, for which it suffices to show the same for $\operatorname{Pic}^{+}_{X/K}(K_{s})=\operatorname{Pic}(X_{K_{s}})$. Because $\operatorname{Pic}(X_{L})=\operatorname{Pic}(\mathbf{A}^{n}_{L})=0$ for some finite extension $L/K$, the claim follows from [Ros4, Lem. 3]. ∎
@@
where $d\colon A\rightarrow\Omega^{1}_{A}$ is the universal derivation, $C^{-1}\colon\Omega^{1}_{A}\rightarrow\Omega^{1}/dA$ is the inverse Cartier map defined by $C^{-1}(fdg)=f^{p}g^{p-1}dg$ (and extended by linearity) and $i\colon\Omega^{1}_{A}\rightarrow\Omega^{1}_{A}/dA$ is the projection. By assumption, ${\rm{H}}^{2}(K,{\mathbf{G}}_{\rm{m}})[p]\neq 0$, so it follows that there exist $\lambda,t\in K$ such that $\phi_{k}(\lambda dt)\neq 0\in\operatorname{Br}(K)$. But we will construct a $K$-form $X$ of $\mathbf{A}^{1}$ such that $\phi_{K}(\lambda dt)$ pulls back to $0\in{\rm{H}}^{2}(X,{\mathbf{G}}_{\rm{m}})$.
@@
$X:=\{Y^{p}+\lambda=t^{p-1}X^{p}-X\}\subset\mathbf{A}^{2}_{K}.$
@@
${X_{K(t^{1/p})}}$${\mathbf{A}^{1}_{K(t^{1/p})}}$${(X,Y)}$${Y-t^{(p-1)/p}X}$${(-S^{p}-\lambda,S-t^{(p-1)/p}S^{p}-t^{(p-1)/p}\lambda)}$${S}$

### v4

@@ 6 Structure of the restricted Picard functor
We may assume that $K=K_{s}$ and that $X$ is nonempty and connected. Because $X$ has dense smooth locus, $X(K)\neq\emptyset$, so Proposition 6.10 implies that $\operatorname{Pic}(X\times S)\twoheadrightarrow\operatorname{Pic}^{+}_{X/K}(S)$ is surjective for every smooth $S/K$. We need to show that there are no nonconstant maps to $\operatorname{Pic}^{+}_{X/K}$ for $S$ unirational. We may assume that $S$ is open in $\mathbf{P}^{1}$. Then every element of $\operatorname{Pic}^{+}_{X/K}(S)$ comes from $\operatorname{Pic}(X\times S)=\operatorname{Pic}(X)$ [EGA, IV4Cor. 21.4.11]. Thus, any map $S\rightarrow\operatorname{Pic}^{+}_{X/K}$ is constant. ∎
@@
The second statement of the theorem follows from Proposition 6.10, because $\mathbf{A}^{n}$ has no non-constant units over a field. Because $\mathbf{A}^{n}$ has no non-constant map to an abelian variety, Theorem 6.9 implies that $\operatorname{Pic}^{+}_{X/K}$ is represented by a smooth $K$-group scheme with finitely-generated component group. Proposition 6.11 implies that $(\operatorname{Pic}^{+}_{X/K})^{0}$ is strongly wound, so it only remains to show that the component group of $\operatorname{Pic}^{+}_{X/K}$ is $p$-power torsion, for which it suffices to show the same for $\operatorname{Pic}^{+}_{X/K}(K_{s})=\operatorname{Pic}(X_{K_{s}})$. For this we may assume that $K=K_{s}$. We have that $\operatorname{Pic}(X_{L})=\operatorname{Pic}(\mathbf{A}^{n}_{L})=0$ for some finite extension $L/K$, so the assertion follows from [Ros4, Lem. 3]. ∎
@@
where $d\colon A\rightarrow\Omega^{1}_{A}$ is the universal derivation, $C^{-1}\colon\Omega^{1}_{A}\rightarrow\Omega^{1}/dA$ is the inverse Cartier map defined by $C^{-1}(fdg)=f^{p}g^{p-1}dg$ (and extended by linearity) and $i\colon\Omega^{1}_{A}\rightarrow\Omega^{1}_{A}/dA$ is the projection. By assumption, ${\rm{H}}^{2}(K,{\mathbf{G}}_{\rm{m}})[p]\neq 0$, so it follows that there exist $\lambda,t\in K$ such that $\phi_{k}(\lambda dt)\neq 0\in\operatorname{Br}(K)$. We will construct a $K$-form $X$ of $\mathbf{A}^{1}$ such that $\phi_{K}(\lambda dt)$ pulls back to $0\in{\rm{H}}^{2}(X,{\mathbf{G}}_{\rm{m}})$.
@@
$X:=\{Y^{p}+\lambda=t^{p-1}Z^{p}-Z\}\subset\mathbf{A}^{2}_{K}.$
@@
${X_{K(t^{1/p})}}$${\mathbf{A}^{1}_{K(t^{1/p})}}$${(Y,Z)}$${Y-t^{(p-1)/p}Z}$${(S-t^{(p-1)/p}S^{p}-t^{(p-1)/p}\lambda,-S^{p}-\lambda)}$${S}$
@@
Example 6.14.
Let $K$ be an imperfect field, and let $U$ be a smooth connected unipotent $K$-group with $\operatorname{Pic}^{+}_{U/K}$ of strictly positive dimension. (Equivalently, since $\operatorname{Pic}^{+}_{U/K}$ is of finite type by Theorem 6.12, $\operatorname{Pic}(U_{K_{s}})$ is infinite.) Such examples exist over every imperfect field; see [KMT, Th. 6.7.9] and [Ros2, Props. 5.2, 5.4, 5.5, Ex. 5.8]. There is a finite étale closed $K$-subgroup scheme $E\subset\operatorname{Pic}^{+}_{U/K}$ such that $\operatorname{Pic}^{+}_{U/K}/E$ is (nontrivial) split unipotent. Indeed, [Ros1, Lem. 2.1.5] implies this for the identity component of $\operatorname{Pic}^{+}_{U/K}$, and we can then further quotient by a finite étale $K$-subgroup of the torsion group $\operatorname{Pic}^{+}_{U/K}$ that surjects onto its component group.
By Theorem 6.12, the inclusion $E\subset\operatorname{Pic}^{+}_{U/K}$ comes from a line bundle $\mathscr{L}\in\operatorname{Pic}(U\times E)$. Let $F/K$ be a finite Galois extension splitting $E$. For each point $e\in E(F)$, $\mathscr{L}_{e}:=\mathscr{L}|U\times\{e\}$ is represented by an effective divisor $D_{e}$ on $U$. Let $D$ be the sum of the Galois conjugates of $\sum_{e\in E(F)}D_{e}$. Then $D$ is a divisor on $U$ such that $\mathscr{L}|(U-D)\times E$ is trivial. Therefore, the composition $E\subset\operatorname{Pic}^{+}_{U/K}\rightarrow\operatorname{Pic}^{+}_{(U-D)/K}$ vanishes. We claim that the map $\operatorname{Pic}^{+}_{U/K}\rightarrow\operatorname{Pic}^{+}_{(U-D)/K}$ is surjective with kernel of strictly smaller dimension than ${\rm{dim}}(\operatorname{Pic}^{+}_{U/K})$ (in fact, the kernel is finite, but we will not require this), from which it will follow that $\operatorname{Pic}^{+}_{(U-D)/K}$ is a nontrivial quotient of the split unipotent $K$-group $\operatorname{Pic}^{+}_{U/K}/E$. We will have therefore constructed over every imperfect field $K$ a smooth connected $K$-scheme whose restricted Picard scheme is nontrivial split unipotent.
To prove the desired surjectivity, it suffices because $\operatorname{Pic}^{+}_{(U-D)/K}$ is smooth to prove surjectivity on $K_{s}$-points. Because all $K_{s}$-points on both schemes come from line bundles, this follows from the surjectivity of $\operatorname{Pic}(U_{K_{s}})\rightarrow\operatorname{Pic}((U-D)_{K_{s}})$, which holds because both schemes are regular and the latter is open in the former. As for the claim about the kernel, if the kernel had dimension equal to that of $\operatorname{Pic}^{+}_{U/K}$, then it would contain all of $(\operatorname{Pic}^{+}_{U/K})^{0}$ and therefore have infinitely many $K_{s}$-points. But the kernel of $\operatorname{Pic}(U_{K_{s}})\rightarrow\operatorname{Pic}((U-D)_{K_{s}})$ is generated by the line bundles associated to the finitely many closed points of $(U-D)_{K_{s}}$. Since it is also torsion, because $\operatorname{Pic}(U_{K_{s}})$ is, it must be finite.
As a consequence of Theorem 6.12, we will deduce that, over a separably closed field, infinitude of the Picard group of a unipotent group is equivalent to infinitude of the translation-invariant line bundles. For this we require the following lemma.
Lemma 6.15.
Let $U,V$ be smooth unipotent $K$-groups equipped with a $U$-action on $V$ ((((as an algebraic group)))).
  • (i)
    If $V\neq 1$, then there is a nontrivial smooth $K$-subgroup of $V$ on which $U$ acts trivially.
  • (ii)
    If ${\rm{dim}}(V)>0$, then there is a strictly positive-dimensional smooth $K$-subgroup of $V$ on which $U$ acts trivially.
Remark 6.16.
For our purposes, we could actually make do with (ii) in the case in which $U$ and $V$ are connected, which would simplify the proof below.
Proof.
Consider the smooth unipotent $K$-group $W:=U\ltimes V$. Define a chain of subgroups $V_{i}$ of $V$ by the formulas
$V_{0}:=V$
$V_{i+1}:=[U,V_{i}],\hskip 7.22743pti\geq 0.$
One readily verifies inductively that the $V_{i}$ are normalized by $U$, and that $V_{i+1}\subset V_{i}$, so in particular, $V_{i}\subset V_{0}=V$. Further, we claim that the $V_{i}$ are smooth. We proceed by induction, the case $i=0$ being immediate. The formation of $V_{i}$ commutes with arbitrary base change, so we may assume that $K=\overline{K}$ to prove smoothness. But $V_{i+1}$ is the image of a map $(U\times V_{i})^{n}\rightarrow W$, hence reduced. Since $K=\overline{K}$, it follows that $V_{i+1}$ is geometrically reduced, hence smooth, as claimed. Now, by definition, $V_{i}\subset\mathscr{D}_{i}(W)$, the $i$th step in the descending central series of $W$. Because $W$ is unipotent, it is nilpotent, hence $\mathscr{D}_{i}(W)=1$ for large $i$. If we then let $i_{0}\geq 0$ be the largest integer such that $V_{i_{0}}\neq 1$, then $V_{i_{0}}$ is centralized by $U$, which proves (i).
For (ii), we first note that $V^{0}$ is normalized by $U$, hence we may assume that $V$ is connected. Now consider the $K$-subgroups $V^{\prime}_{i}$ of $V$ defined as follows:
$V^{\prime}_{0}:=V$
$V^{\prime}_{i+1}:=[U^{0},V^{\prime}_{i}],\hskip 7.22743pti\geq 0.$
Then one readily verifies inductively that $V^{\prime}_{i}$ is normalized by $U$ and contained in $V^{\prime}_{i+1}$. It is connected by a similar argument to that given for smoothness of the $V_{i}$ above. Just as with the $V_{i}$ we conclude that $V^{\prime}_{i}=1$ for all large enough $i$. If we let $i_{0}\geq 0$ be the maximal integer such that $V^{\prime}_{i_{0}}\neq 1$, then $V^{\prime}_{i_{0}}\subset V$ is smooth, connected, positive-dimensional, normalized by $U$, and acted on trivially by $U^{0}$. Replacing $V$ by $V^{\prime}_{i_{0}}$, therefore, and $U$ by $U/U^{0}$, we may assume that $U=E$ is finite étale.
The $E$-invariants $V^{E}$ of $V$ form a closed $K$-subgroup scheme of $V$. The assertion we wish to prove is that the maximal smooth $K$-subgroup scheme of $V^{E}$ is positive-dimensional. Since the formation of both $V^{E}$ and the maximal smooth subgroup scheme commute with separable base change [CGP, Lem. C.4.1], we are free to assume that $K$ is separably closed. We then prove the assertion by induction on #$\#E$, the case #$\#E=1$ being trivial. So suppose that $E\neq 1$ is an étale unipotent $K$-group. Because $E$ is unipotent, it has nontrivial center, so choose $1\neq e\in E(K)$ lying in the center of $E$.
Consider as before the $K$-group $W:=E\ltimes V$, and define the $K$-subgroups $V^{\prime\prime}_{i}\subset V$ by the following formula:
$V^{\prime\prime}_{0}:=V$
[…]
where $[e,V^{\prime\prime}_{i}]$ means the group generated by the map $V^{\prime\prime}_{i}\rightarrow W,v\mapsto[e,v]$. As before, one readily verifies that $V^{\prime\prime}_{i}$ is a descending series of smooth connected subgroups of $V$, all of which are normalized by $E$ (this uses the fact that $e$ is central in $E$), and again, as before, unipotence of $W$ implies that there is some $i_{0}$ such that $V^{\prime\prime}_{i_{0}}$ is positive-dimensional and $e$ acts trivially on $V^{\prime\prime}_{i_{0}}$. Replacing $V$ by $V^{\prime\prime}_{i_{0}}$, then, we may suppose that $e$ acts trivially on $V$. Then the action descends to an action of $E/\langle e\rangle$ on $V$, and we are therefore done by induction. ∎
Recall that, for a smooth connected $K$-group scheme $G$,
$\operatorname{Ext}^{1}(G,{\mathbf{G}}_{\rm{m}}):=\{\mathscr{L}\in\operatorname{Pic}(G)\mid m^{*}\mathscr{L}\simeq\pi_{1}^{*}\mathscr{L}\otimes\pi_{2}^{*}\mathscr{L}\}\subset\operatorname{Pic}(G).$
That is, $\operatorname{Ext}^{1}(G,{\mathbf{G}}_{\rm{m}})$ consists of the universally translation-invariant line bundles on $G$. For an explanation of the notation, see the introduction to [Ros2].
Proposition 6.17.
Let $U$ be a smooth connected unipotent $K$-group.
  • (i)
    $\operatorname{Pic}(U_{K_{s}})$ is trivial if and only if $\operatorname{Ext}^{1}_{K_{s}}(U,{\mathbf{G}}_{\rm{m}})$ is.
  • (ii)
    $\operatorname{Pic}(U_{K_{s}})$ is finite if and only if $\operatorname{Ext}^{1}_{K_{s}}(U,{\mathbf{G}}_{\rm{m}})$ is.
Proof.
We may assume that $K$ is separably closed. The only if directions are immediate, so we only need to prove the if directions. By Theorem 6.12, $\operatorname{Pic}^{+}_{U/K}$ is represented by a (smooth) unipotent $K$-group scheme. Furthermore, $\operatorname{Pic}^{+}_{U/K}(K)=\operatorname{Pic}(U)$. Now the action of $U$ on itself by left-translation induces an action on $\operatorname{Pic}^{+}_{U/K}$, and $x\in\operatorname{Pic}^{+}_{U/K}(K)$ is fixed by $U$ if and only if the corresponding line bundle lies in $\operatorname{Ext}^{1}(U,{\mathbf{G}}_{\rm{m}})$. Thanks to Lemma 6.15, we have that $((\operatorname{Pic}^{+}_{U/K})^{U})^{\rm{sm}}$ (with ${\rm{sm}}$ denoting the maximal smooth $K$-subgroup scheme) is trivial, respectively étale, if and only if the same holds for $\operatorname{Pic}^{+}_{U/K}$, which proves the proposition. ∎